Logiweb bibliography of finalversion

Up Help

[0] finalversion

[1] base

[2] check

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-07-14.UTC:09:09:53.060274 = MJD-53930.TAI:09:10:26.060274 = LGT-4659585026060274e-6